# DAS_CE_Supp

### v2

@@ IV-D Majorization Minimization for Phase-Only Controllable Combiner
It can be easily proved that the optimal solution to (IV-D) is given by ${{\bf{w}}}_{t+1}=\frac{1}{{\sqrt{M}}}{\exp\left(j\angle\left({{\bf{X}}-{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}+{{\bf{\Sigma}}_{t}}}\right){\bf{v}}\right)}$. This solution, associated with the updating approach of the posterior kernel in Lemma 1, completes the algorithm. It is worth noting that, since the updates of $\bf v$ and ${\bf w}_{t+1}$ both monotonously increase the objective function, the convergence of Algorithm 2 is naturally guaranteed.
@@ IV-E Kernel Selection in Non-Ideal Case
In practical communication scenarios, obtaining the perfect prior kernel, ${\bf\Sigma}_{\bf h}$, in the initialization states of Algorithms 1 and 2 may pose a challenge. To deal with this issue, we suggest two methods to obtain imperfect prior kernels.
@@ IV-E2 Artificial kernels
    In comparison to other kernels, the exponential kernel demonstrates lower sensitivity to outliers, rendering it suitable for reconstructing channels lacking obvious regularity.
  • Bessel kernel: The Bessel kernel, denoted as ${\bm{\Sigma}}_{\rm bes}$, is well-suited for capturing and modeling complex-valued data exhibiting oscillatory patterns. The Bessel kernels for the two dimensions of UPA can be expressed as
@@
    ${\bm{\Sigma}}_{\rm bes}$ possesses the flexibility to adapt to data that features regular and repeats fluctuations, which might not be adequately modeled by alternative kernels.
@@ V-A Estimation Accuracy Under Perfect Kernel
It is evident from Lemmas 2-4 that meticulously designed observation matrices can scale down the asymptotic MSE from $\mathcal{O}(KMQ^{-1})$ to $\mathcal{O}(K^{2}Q^{-1})$. This improvement is attributed to the fact that the observation matrices designed by water-filling and ice-filling approaches are tightly aligned with the kernel’s eigenspace of dimension $K$. In contrast, the random observation matrix wastes a large amount of power in the kernel’s null space of dimension $M-K$, which contains no useful information about the channel. This finding demonstrates the significant role played by the observation matrix to enhance channel estimation accuracy, particularly in DASs where $M\gg K$.
@@ V-C Computational Complexity Analysis
For the observation matrix design, the complexity of Algorithm 1 is mainly caused by the eigenvalue decomposition and the principal eigenvalue selection, which is ${\cal O}(M^{3}+QK)$. The complexity of Algorithm 2 mainly depends on the search for the largest eigenvalue and the iterations required for the convergence of MM, which is ${\cal O}(M^{3}+I_{o}M^{2}Q)$ wherein $I_{o}$ is the number of iterations. and then saved in the memory medium in advance.
@@ VI-A Simulation Setup and Benchmarks
To account for a practical scenario, the standard 3GPP TR 38.901 channel model is used for simulations, whose key parameters are given in Table I. Otherwise specifically specified, we consider a UPA. The number of antennas is set to $M=128$, and the numbers of horizontal antennas and vertical antennas are set to $M_{x}=16$ and $M_{y}=8$, respectively. Assume that the antenna spacing is set to $\frac{\lambda}{8}$. The signal-to-noise ratio (SNR) is defined as ${\rm SNR}=\frac{{\mathsf{E}}\left(\|{\bf h}\|^{2}\right)}{\sigma^{2}}$, of which the default value is set to $10$ dB. The performance is evaluated by the normalized mean square error (NMSE), which is defined as ${\rm NMSE}={\mathsf{E}}\left(\frac{\|{\bf h}-\hat{\bf h}\|^{2}}{\|{\bf h}\|^{2}}\right)$. Besides, 3000 Monte-Carlo experiments are carried out to plot each figure.
To demonstrate the effectiveness of the proposed ice-filling and MM enabled channel estimators, five benchmark schemes are considered for comparison:
  • LS: The classic LS method is applicable when the pilot length is no less than the number of antennas. To implement it, the pilot length is fixed as $Q=M$ and the observation matrix, ${\mathbf{W}}$, is set to a DFT matrix.
  • MMSE: The MMSE benchmark utilizes the same DFT observation matrix as the LS scheme. It then leverages the Bayesian regression in (4) to recover the channel.
  • AMP: The vectored approximate message passing (AMP) method proposed in [31] is leveraged to estimate the sparse channel in angular domain.
  • Ideal water-filling: The water-filling scheme is employed to design the ideal observation matrix ${\bf W}$. The noise vector, ${\mathbf{z}}$, in (2) is ideally assumed to be independent of the observation matrix ${\mathbf{W}}$, and its distribution is fixed as $\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$.
  • Random W: An observation matrix ${\mathbf{W}}$ of $Q$-pilot length is randomly generated with standard complex Gaussian distribution and utilized for Bayesian regression.
Note that the pilot length for the LS and MMSE methods is $Q=M=128$, while that for the other schemes is set to $Q=64$ by default.
@@ VI-B Performance Evaluation Under Perfect Kernel
In this subsection, we use the perfect kernel, ${\bm{\Sigma}}_{{\mathbf{h}}}$, to trigger the water-filling, ice-filling, and MM algorithms, and then perform the Bayesian regression. The NMSE performance as a function of SNR is evaluated in Fig. 4. The SNR ranges from $-10{\>}{\text{dB}}\sim 15{\>}{\text{dB}}$. Thanks to the careful design of observation matrix, the ideal water-filling, the proposed ice-filling, and the proposed MM algorithms remarkably outperform existing benchmarks. In particular, a 5 dB NMSE gap between ice-filling and MMSE, as well as a 15 dB gap between ice-filling and AMP, are visible when the SNR is 10 dB. Notably, the NMSE performance attained by the ice-filling method tightly aligns with that of the ideal water-filling scheme, given that the ice-filling intrinsically allocates $Q$ quantized units of power to the $Q$-timeslot pilots. Another finding of interest is that the NMSE gap between the MM and ice-filling algorithms is no higher than 1 dB, demonstrating the near-optimality of the proposed MM algorithm in phase-only-controllable design.
Figure 5: The NMSE as a function of pilot length. The pilot length allocated to the LS and MMSE methods is fixed to 128, whereas the pilot length $Q$ for other schemes rises from 8 to 68.
@@
[…]
In Fig. 6, we demonstrate the superiority of a DAS in precise channel estimation. The curves of the achieved NMSE performance versus the ratio of antenna spacing and wavelength, i.e., $\frac{d}{\lambda}$, are plotted. We can observe that, as the normalized antenna spacing $\frac{d}{\lambda}$ decreases from $\frac{1}{2}$ to $\frac{1}{16}$, the NMSEs for all channel estimators excluding LS method declines rapidly. For example, the NMSE achieved by the proposed ice-filling algorithm is decreased by about 10 dB when the antenna spacing ranges from $\lambda/2$ to $\lambda/8$. This phenomenon is attributed to the fact that, a smaller antenna spacing leads to stronger spatial correlations of channels over different antennas, which results in a more informative kernel ${\bm{\Sigma}}_{{\mathbf{h}}}$ for more precise channel reconstruction. Since the underpinning spatial information of wireless channels is embedded in the structured kernel, these Bayesian optimization based channel estimators are enhanced under DASs.
@@ VI-C Performance Evaluation Under Imperfect Kernel
In some scenarios where the perfect prior kernel $\bm{\Sigma}_{\bf h}$ cannot be obtained, the statistical kernel and artificial kernels defined in Section IV-E can be utilized to replace the perfect kernel $\bm{\Sigma}_{\bf h}$ in Algorithms 1$\sim$2. We first consider the statistical kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}={\bf{\Sigma}}_{{\bf h}}+\sigma^{2}_{{\mathbf{h}}}{\mathbf{I}}_{M}$. The NMSE as a function of the kernel estimation error $\sigma^{2}_{{\mathbf{h}}}$ is shown in Fig. 7. One can observe that all kernel-dependent methods, including the MMSE, Random ${\mathbf{W}}$, and the proposed estimators, suffer from visible performance loss due to the imperfect kernel input. Fortunately, the proposed methods still hold the superiority among all schemes. In particular, the estimation accuracy of the proposed methods is almost not influenced when $\sigma^{2}_{{\mathbf{h}}}<-15\,{\text{dB}}$, which shows their high robustness to the kernel error. One more interesting finding is that, when $\sigma^{2}_{{\mathbf{h}}}$ increases, the gap between MMSE and our proposed methods, become small. The reason is that, an increasing $\sigma^{2}_{{\mathbf{h}}}$ gradually obscures the structural characteristics of ${\bf{\Sigma}}_{{\bf h}}$, which makes the input kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}$ behaves more like an identity matrix. In this case, the importance of designing the observation matrix $\bf W$ is reduced.
Figure 8: The effect of SNR on NMSE performance under the perfect kernel ${\bm{\Sigma}}_{\bf h}$, the Bessel kernel ${\bm{\Sigma}}_{\rm bes}$, and the exponential kernel ${\bm{\Sigma}}_{\rm exp}$.
We further evaluate the achievable NMSE performance under the artificial kernels $\bm{\Sigma}_{\rm exp}$ and ${\bm{\Sigma}}_{\rm Bes}$. Their hyper-parameters, $\eta_{1}$ and $\eta_{2}$, are determined by the binary search through several numerical experiments, which are set to $\eta_{1}=0.56$ and $\eta_{2}=0.85$ in the considered scenario. In this way, we obtain the NMSE versus the SNR and that versus the pilot length $Q$ in Fig. 8 and Fig. 9, respectively. From these two figures one find that, for a given kernel, the curves of “ideal water-filling”, “proposed ice-filling”, and “proposed MM” are very close, which indicates that these three schemes have similar robustness to imperfect kernels. This phenomenon proves the feasibility of using experiential kernels.
@@ VII Conclusions
This paper incorporated the design of observation matrix into Bayesian channel estimation in DASs. The formulated MIM-based observation matrix design was shown to be a time-domain duality of classic MIMO precoding, which was ideally addressed by the water-filling principle. Targeting practical DAS realizations, we proposed a novel ice-filling and a MM enabled algorithms to design amplitude-and-phase controllable and phase-only controllable observation matrices, respectively. In particular, the ice-filling algorithm was proved to be a discrete approximation of water-filling, with the relative quantization error decaying rapidly. Comprehensive analyses on the MSEs of channel estimation and their asymptotic expressions validated the near-optimality of the proposed designs. Last, numerical results illustrated that our algorithms could improve the channel estimation accuracy by orders of magnitude.
@@
Appendix A Proof of Lemma 1
The key to this proof lies in calculating the inverse of ${{{\bf{W}}_{t}^{H}}{{{\bf{\Sigma}}_{\bf{h}}}{{{{{\bf{W}}_{t}}}}}+{\sigma^{2}}{{\bf{I}}_{t}}}}$ in (9). For ease of notation, we denote ${\bf{K}}_{t}={{{\bf{W}}_{t}^{H}}{{{\bf{\Sigma}}_{\bf{h}}}{{{{{\bf{W}}_{t}}}}}+{\sigma^{2}}{{\bf{I}}_{t}}}}$ and ${\bf{W}}_{t}=[{{\bf{W}}_{t-1}},{\bf{w}}_{t}]$. Then ${\bf{K}}_{t}$ can be expressed in a block matrix form:

### v3

@@ IV-D Majorization Minimization for Phase-Only Controllable Combiner
It can be easily proved that the optimal solution to (IV-D) is given by ${{\bf{w}}}_{t+1}=\frac{1}{{\sqrt{M}}}{\exp\left(j\angle\left({{\bf{X}}-{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}+{{\bf{\Sigma}}_{t}}}\right){\bf{v}}\right)}$. This solution, associated with the updating approach of the posterior kernel in (9), completes the algorithm. It is worth noting that, since the updates of $\bf v$ and ${\bf w}_{t+1}$ both monotonously increase the objective function, the convergence of Algorithm 2 is naturally guaranteed.
@@ IV-E Kernel Selection in Non-Ideal Case
In some non-ideal cases where the perfect prior kernel, ${\bf\Sigma}_{\bf h}$, is difficult to acquire, the initialization states of Algorithms 1 and 2 may pose a challenge. To deal with this issue, we suggest two methods to obtain imperfect prior kernels.
@@ IV-E2 Artificial kernels
    The exponential kernel, ${\bm{\Sigma}}_{\rm exp}$, is a heuristic kernel. It exhibits the highest correlation at the diagonal elements, with the channel correlation diminishing exponentially as the antenna spacing increases. Compared to other kernels, the exponential kernel shows reduced sensitivity to outliers, making it suitable for reconstructing channels that lack obvious regularity.
  • Bessel kernel: The Bessel kernel, denoted as ${\bm{\Sigma}}_{\rm bes}$, is well-suited for capturing and modeling complex-valued data exhibiting oscillatory patterns. It characterizes the covariance of array steering vector when the user distributes uniformly in the spatial domain. For the considered UPA, the Bessel kernel can be obtained as ${\bm{\Sigma}}_{\rm bes}=\mathsf{E}({\mathbf{a}}(\theta,\phi){\mathbf{a}}^{H}(\theta,\phi))$, where ${\mathbf{a}}(\theta,\phi)$ is the array steering vector of UPA, and $\theta\sim\mathcal{U}(-\pi/2,\pi/2)$ and $\phi\sim\mathcal{U}(-\pi/2,\pi/2)$ denote the angle-of-arrivals. By calculating expectation, the Bessel kernel is expressed as
@@ V-A Estimation Accuracy Under Perfect Kernel
It is evident from Lemmas 1-3 that meticulously designed observation matrices can scale down the asymptotic MSE from $\mathcal{O}(KMQ^{-1})$ to $\mathcal{O}(K^{2}Q^{-1})$. This improvement is attributed to the fact that the observation matrices designed by water-filling and ice-filling approaches are tightly aligned with the kernel’s eigenspace of dimension $K$. In contrast, the random observation matrix wastes a large amount of power in the kernel’s null space of dimension $M-K$, which contains no useful information about the channel. This finding demonstrates the significant role played by the observation matrix to enhance channel estimation accuracy, particularly in DASs where $M\gg K$.
@@ V-C Computational Complexity Analysis
For the observation matrix design, the complexity of Algorithm 1 is mainly caused by the eigenvalue decomposition and the principal eigenvalue selection, which is ${\cal O}(M^{3}+QK)$. The complexity of Algorithm 2 mainly depends on the search for the largest eigenvalue and the iterations required for the convergence of MM, which is ${\cal O}(M^{3}+I_{o}M^{2}Q)$ wherein $I_{o}$ is the number of iterations.
@@ VI-A Simulation Setup and Benchmarks
To account for a practical scenario, the 3GPP TR 38.901 channel model is used for simulations, whose key parameters are given in Table I. Otherwise specifically specified, we consider a UPA. The number of antennas is set to $M=128$, and the numbers of horizontal antennas and vertical antennas are set to $M_{x}=16$ and $M_{y}=8$, respectively. The antenna spacing is set to $\frac{\lambda}{8}$. The signal-to-noise ratio (SNR) is defined as ${\rm SNR}=\frac{{\mathsf{E}}\left(\|{\bf h}\|^{2}\right)}{\sigma^{2}}$. The performance is evaluated by the normalized mean square error (NMSE), which is defined as ${\rm NMSE}={\mathsf{E}}\left(\frac{\|{\bf h}-\hat{\bf h}\|^{2}}{\|{\bf h}\|^{2}}\right)$. 3000 Monte-Carlo experiments are carried out to plot each figure.
To demonstrate the effectiveness of the proposed ice-filling and MM algorithms, six observation matrices are considered for comparison:
  • Random matrix ${\mathbf{W}}^{\rm Rand}\in\mathbb{C}^{M\times Q}$: The coefficients of ${\mathbf{W}}^{\rm Rand}$ are independently generated from the standard complex Gaussian distribution. Then, each column is normalized to satisfy the unit power constraint.
  • Top-$Q$ matrix ${\mathbf{W}}^{{\rm Top}Q}\in\mathbb{C}^{M\times Q}$: The first $Q$ major eigenvectors of ${\bf\Sigma}_{{\mathbf{h}}}$ are used to set the observation matrix ${\mathbf{W}}^{{\rm Top}Q}$.
  • Ice-filling matrix ${\mathbf{W}}^{\rm IF}\in\mathbb{C}^{M\times Q}$: The proposed ice-filling algorithm is used to generate ${\mathbf{W}}^{\rm IF}$.
  • MM matrix ${\mathbf{W}}^{\rm MM}\in\mathbb{C}^{M\times Q}$: The proposed MM algorithm is used to generate ${\mathbf{W}}^{\rm MM}$.
  • Water-filling matrix ${\mathbf{W}}^{\rm WF}\in\mathbb{C}^{M\times Q}$: The water-filling scheme is employed to generate ${\mathbf{W}}^{\rm WF}$. To implement this ideal observation matrix, the noise vector, ${\mathbf{z}}$, in (2) is assumed to be independent of the observation matrix ${\mathbf{W}}^{\rm WF}$, and its distribution is fixed as $\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$.
  • DFT matrix ${\mathbf{W}}^{\rm DFT}\in\mathbb{C}^{M\times M}$: The DFT matrix requires the pilot length, $Q$, to be equal to the number of antennas, $M$. This matrix is used for performing LS algorithm.
Note that the pilot length for the DFT matrix ${\mathbf{W}}^{\rm DFT}$ is $Q=M=128$, while that for the other matrices is $Q=64$ by default. The above observation matrices are used in four channel estimators for comparison: The LS, vector approximate message passing (VAMP) [28], OMP, and MMSE algorithms. For the LS algorithm, it has to be implemented by the DFT matrix ${\mathbf{W}}^{\rm DFT}$ with a $Q=M$ pilot length. As for the OMP and VAMP algorithms, we conduct numerical experiments in advance to evaluate their performance under matrices, ${\mathbf{W}}^{\rm Rand}$, ${\mathbf{W}}^{{\rm Top}Q}$, ${\mathbf{W}}^{\rm MM}$ and ${\mathbf{W}}^{\rm IF}$, respectively. It is shown that the combinations of “VAMP + ${\mathbf{W}}^{\rm Rand}$” and “OMP + ${\mathbf{W}}^{{\rm Top}Q}$” exhibit the best channel estimation performance [^2] This is because CS algorithms expect to make the column coherence of observation matrix as small as possible. As the designed matrix ${\mathbf{W}}^{{\rm IF}}$ might select the same eigenvectors in different time slots, it is not suitable for OMP and VAMP algorithms.. Thereby, the following simulations use the random matrix ${\mathbf{W}}^{\rm Rand}$ for VAMP and the top-$Q$ matrix ${\mathbf{W}}^{{\rm Top}Q}$ for OMP. The MMSE algorithm in (4) is the estimator adopted in this paper. We will evaluate its performance under matrices, ${\mathbf{W}}^{\rm Rand}$, ${\mathbf{W}}^{{\rm Top}Q}$, ${\mathbf{W}}^{\rm MM}$, ${\mathbf{W}}^{\rm IF}$, and ${\mathbf{W}}^{\rm WF}$, respectively, to show the superiority of the designed observation matrices.
@@ VI-B Performance Evaluation Under Perfect Kernel
In this subsection, we use the perfect kernel, ${\bm{\Sigma}}_{{\mathbf{h}}}$, to trigger the water-filling, ice-filling, and MM algorithms, and then perform the Bayesian regression. The NMSE performance as a function of SNR is evaluated in Fig. 4. The SNR ranges from $-10{\>}{\text{dB}}\sim 10{\>}{\text{dB}}$. Thanks to the careful design of observation matrix, our schemes “MMSE + ${\mathbf{W}}^{\rm MM}$” and “MMSE + ${\mathbf{W}}^{\rm IF}$” remarkably outperform existing benchmarks. For example, a 10 dB NMSE gap between “MMSE + ${\mathbf{W}}^{\rm IF}$” and “VAMP + ${\mathbf{W}}^{\rm Rand}$ is visible. Moreover, the ice-filling matrix ${\mathbf{W}}^{\rm IF}$ can reduce the NMSE of MMSE algorithms by 5 dB compared to the top-$Q$ matrix ${\mathbf{W}}^{{\rm Top}Q}$. Notably, the NMSE performance of MMSE attained by the ice-filling matrix tightly aligns with that of the ideal water-filling matrix, given that the ice-filling intrinsically allocates $Q$ quantized units of power to the $Q$-timeslot pilots. Another finding of interest is that the NMSE gap between “MMSE + ${\mathbf{W}}^{\rm IF}$” and “MMSE + ${\mathbf{W}}^{\rm MM}$” is no higher than 0.5 dB, demonstrating the near-optimality of the proposed MM algorithm in phase-only-controllable design.
Figure 5: The NMSE as a function of pilot length. The pilot length allocated to the LS method is fixed to 128, whereas the pilot length $Q$ for other schemes rises from 8 to 68.
@@
Next, we investigate the impact of pilot length, $Q$, on the NMSE performance in Fig. 5. The pilot length allocated to the LS method is fixed to 128, whereas the pilot length for the other algorithms ranges from 8 to 68. The SNR is -5 dB. It is evident from Fig. 5 that the proposed algorithms consume significantly lower pilot overhead than the other benchmarks to achieve the same NMSE level. For instance, to obtain a $-5\,\text{dB}$ NMSE, the pilot length required for the VAMP algorithm is larger than 30. In contrast, 8 pilots are sufficient for the proposed ice-filling and MM algorithms to attain a NMSE lower than - 10 dB. Additionally, it is notable that the NMSE performance of “MMSE + ${\mathbf{W}}^{{\rm Top}Q}$ no longer decreases when $Q>12$. This is because the top-$Q$ matrix begins to allocate pilots to invalid eigenvectors with very small eigenvalues when $Q>12$, while the proposed ice-filling and MM algorithms continue to select valid eigenvectors from the $1^{\rm st}$ to $12^{\rm th}$ eigenvalues. As a result, the performance gain attained by our algorithms continuously improve as the pilot length increases.
In Fig. 6, we demonstrate the superiority of a DAS in precise channel estimation. The curves of the achieved NMSE performance versus the ratio of antenna spacing and wavelength, i.e., $\frac{d}{\lambda}$, are plotted. The SNR is 0 dB. We can observe that, as the normalized antenna spacing $\frac{d}{\lambda}$ decreases from $\frac{3}{2}$ to $\frac{1}{16}$, the NMSEs for all channel estimators excluding LS method declines rapidly. For example, the NMSE achieved by the proposed ice-filling algorithm is decreased by about 15 dB when the antenna spacing ranges from $\lambda/2$ to $\lambda/8$. This phenomenon is attributed to the fact that, a smaller antenna spacing leads to stronger spatial correlations of channels over different antennas, which results in a more informative kernel ${\bm{\Sigma}}_{{\mathbf{h}}}$ for more precise channel reconstruction. Since the underpinning spatial information of wireless channels is embedded in the structured kernel, these Bayesian regression based channel estimators are enhanced under DASs.
@@ VI-C Performance Evaluation Under Imperfect Kernel
In some scenarios where the perfect prior kernel $\bm{\Sigma}_{\bf h}$ cannot be obtained, the statistical kernel and artificial kernels defined in Section IV-E can be utilized to replace the perfect kernel $\bm{\Sigma}_{\bf h}$ in Algorithms 1$\sim$2. We first consider the statistical kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}={\bf{\Sigma}}_{{\bf h}}+\sigma^{2}_{{\mathbf{h}}}{\mathbf{I}}_{M}$. The NMSE as a function of the kernel estimation error $\sigma^{2}_{{\mathbf{h}}}$ is shown in Fig. 7. One can observe that all MMSE methods suffer from visible performance loss due to the imperfect kernel input. Fortunately, the proposed methods still hold the superiority among all schemes. In particular, the estimation accuracy of the proposed methods is almost not influenced when $\sigma^{2}_{{\mathbf{h}}}<-15\,{\text{dB}}$, which shows their high robustness to the kernel error.
Figure 8: The effect of SNR on NMSE performance under the perfect kernel ${\bm{\Sigma}}_{\bf h}$, the statistical kernel $\hat{{\bm{\Sigma}}}_{{\mathbf{h}}}$, the Bessel kernel ${\bm{\Sigma}}_{\rm bes}$, and the exponential kernel ${\bm{\Sigma}}_{\rm exp}$.
We further evaluate the achievable NMSE performance under all considered kernels, including the perfect kernel ${\bm{\Sigma}}_{{\mathbf{h}}}$, the statistical kernel $\hat{{\bm{\Sigma}}}_{{\mathbf{h}}}$, and the artificial kernels $\bm{\Sigma}_{\rm exp}$ and ${\bm{\Sigma}}_{\rm bes}$. The estimation error of $\hat{{\bm{\Sigma}}}_{{\mathbf{h}}}$ is set as ${\mathbf{\sigma}}_{{\mathbf{h}}}^{2}=-12\>{\rm dB}$. The hyper-parameters, $\eta_{1}$ and $\eta_{2}$, of $\bm{\Sigma}_{\rm exp}$ and ${\bm{\Sigma}}_{\rm bes}$ are determined by the binary search through several numerical experiments, which are set to $\eta_{1}=0.56$ and $\eta_{2}=0.85$ in the considered scenario. In this way, we obtain the NMSE versus the SNR and that versus the pilot length $Q$ in Fig. 8 and Fig. 9, respectively. For clarity, we only plot the curves of MMSE algorithms. From these two figures one find that, for a given kernel, the curves of “ideal water-filling”, “proposed ice-filling”, and “proposed MM” are very close, which indicates that these three schemes have similar robustness to imperfect kernels.
@@ VII Conclusions
This paper incorporated the design of observation matrix into Bayesian channel estimation in DASs. The formulated MIM-based observation matrix design was shown to be a time-domain duality of classic MIMO precoding, which was ideally addressed by the water-filling principle. Targeting practical DAS realizations, we proposed a novel ice-filling and a MM enabled algorithms to design amplitude-and-phase controllable and phase-only controllable observation matrices, respectively. In particular, the ice-filling algorithm was proved to be a discrete approximation of water-filling, with the relative quantization error decaying rapidly. Comprehensive analyses and numerical results validated the near-optimality of the proposed designs.
@@
Appendix A Proof of (9)
The posterior kernel $\bm{\Sigma}_{t}$ is expressed as
The key to the proof of (9) lies in calculating the inverse of ${{{\bf{W}}_{t}^{H}}{{{\bf{\Sigma}}_{\bf{h}}}{{{{{\bf{W}}_{t}}}}}+{\sigma^{2}}{{\bf{I}}_{t}}}}$ in (45). For ease of notation, we denote ${\bf{K}}_{t}={{{\bf{W}}_{t}^{H}}{{{\bf{\Sigma}}_{\bf{h}}}{{{{{\bf{W}}_{t}}}}}+{\sigma^{2}}{{\bf{I}}_{t}}}}$ and ${\bf{W}}_{t}=[{{\bf{W}}_{t-1}},{\bf{w}}_{t}]$. Then ${\bf{K}}_{t}$ can be expressed in a block matrix form: